can you help me please
given f(x)={x^2 -3 for x > 4
{log(x-3) for x <= 4 , evaluate f(4)

Found 2 solutions by solver91311, ikleyn:
Answer by solver91311(24713)



Since 4 is not greater than 4, ignore the piece of the piecewise function.

The piece of the piecewise function applies because 4 is indeed less than or equal to 4.

Write the applicable piece of the function in pencil and then every place you see an , erase it an write a 4 in its place.

Then do the arithmetic. Hint:


John

My calculator said it, I believe it, that settles it


Answer by ikleyn(52781)

.
can you help me please
given f(x) = {x^2 -3 for x > 4
             {log(x-3) for x <= 4, 

evaluate f(4)
~~~~~~~~~~~~~~~


Hello, it is easy.

    The value of 4  belongs to the set  {x <= 4}  and does not belong to the set {x > 4}.


    Right ?  Do you understand it ?


    But of course ! --- It is OBVIOIUS (!)


    Therefore, you should use the FIRST line of the two that define the function.


    So, substitute x= 4 into the formula in the first line of the definition.


    You will get then  f(x) = log(4-3) = log(1) = 0.      ANSWER


That is all.